# Quantum Hermitian conjugate and encoding unnormalized matrices

### v2

@@ Abstract
We further develop the family of matrix-manipulation algorithms based on the encoding the matrix elements into the probability amplitudes of the pure superposition state of a certain quantum system. We introduce two extensions to these algorithms which allow (i) to perform Hermitian conjugation of matrices under consideration and (ii) to weaken the restriction to the absolute values of matrix elements unavoidably imposed by the normalization condition for a pure quantum state. Both these extensions are applied to the matrix multiplication algorithm. Controlled measurement of ancilla state is implemented to avoid the problem of small success probability in the measurement process.
@@ I Introduction
Quantum algorithm is a rapidly developing area of quantum informatics whose progress is stimulated by prospective enormous advantages of quantum algorithms over their classical counterparts. We have to acknowledge such famous algorithms as the Deutsch algorithm Deu demonstrating effect of quantum parallelism in operation with Boolean functions, Shor’s factoring algorithm Sho1 ; Sho2 utilising the quantum Fourier transform QFT1 ; QFT2 ; QFT3 , Grover’s search algorithm Gro , phase-estimation algorithm for the Hermitian QFT3 and non-Hermitian Wang matrices .
@@
We consider a particular application area of quantum algorithms which covers matrix algebraic operations, such as multiplication, addition, determinant calculation, matrix inversion and its application to linear system solver. During recent years, there are many quantum algorithms solving various aspects of linear algebra. Among them are HHL-algorithm for solving systems of linear equations HHL which was applied, in particular, to Gaussian process regression ZhaoZ , algorithm for solving linear differential equations Berry , algorithm for evaluation of matrix functions Tak .
@@
A possible approach to quantum algorithms for matrix manipulations is proposed in ZZRF . This approach refers to matrices as operators to be applied to some vector. Therefore, it is essentially in that approach that arbitrary nonunitary and non-Hermitian matrix must be transferred to the unitary one. Thus, first one has to replace a given matrix $A$ (rectangular in general) with the Hermitian one and then transfer it to the unitary operator by exponentiating which can be approximated via Trotterization Trotter ; Suzuki ; T ; AT and Baker-Champbell-Hausdorff method BC . To implement such matrix transformations effectively, the set of special matrices is proposed to be prepared in advance. The phase estimation subroutine QFT3 is included into that algorithm.
@@
A special block-encoding model GSLW is another approach for implementing matrix algebra. Different type of block-encoding was used in the algorithm in Ref.DFZ_2020 to embed the inverse of the matrix into the unitary transformation. In LWWZ , matrix multiplication was performed trough the binary encoding of matrix elements into the pure states of quantum systems with subsequent binary multiplication via Toffoli gates. Matrix multiplication over some rings, in particular, multiplication of Boolean matrices, was studied in KN .
@@
Some algorithms for matrix inversion have been also developed. As the first one we call the mentioned above HHL algorithm HHL where the inversion was used for solving a linear system. However, the problem of inversion of matrix eigenvalues has not been resolved in frames of quantum algorithm therein. The matrix inversion algorithm based on the singular value decomposition is presented in MRTC ; NJ . In both paper the authors apply the function evaluation problem GSLW to approximate the inverse of singular values by odd-polynomial. This approach requires introducing a special scale parameter depending on the interval where all nonzero singular values are distributed.
@@
Recently the alternative matrix-manipulation algorithms based on the matrix encoding into the complex probability amplitudes of a pure superposition state of a certain subsystem have been developed QZKW_arxive2022 ; ZQKW_2024 ; ZBQKW_arXive2024 . The principal difference of such approach is that a matrix is considered as a part of the quantum state of a system rather then a unitary operator applied to this system. Therefore the elements of the resulting matrix (obtained after addition or multiplication of other matrices, or after matrix inversion) can be measured up to some constant factor. Although those algorithms deal with complex matrices in general, some operations (including Hermitian conjugation) can not be realized in frames of those settings and therefore certain constraints are to be imposed on the variety of matrix manipulations. To remove this disadvantage and construct more flexible algorithms for matrix algebra we propose to separate the real and imaginary parts of a matrix by encoding them in different probability amplitudes of a pure state. This separation can be done by involving additional one-qubit subsystem whose state serves to label the real and imaginary parts of a matrix elements. As a consequence, all probability amplitudes are real in such representation.
Another issue addressed in this paper is weakening the normalization constraint for the elements of the considered matrices. We recall that such constraint is imposed by the normalization condition for a pure quantum state. For that purpose we introduce an additional probability amplitude into the superposition state encoding considered matrix. This can be done involving one more auxiliary one-qubit subsystem.
@@
Of course, both above extensions require modifications of the algorithms proposed in ZQKW_2024 ; ZBQKW_arXive2024 . Below we present the detailed analysis of the appropriate modifications introduced into the multiplication algorithm (see appendix in ZBQKW_arXive2024 ) and demonstrate manipulations with the input matrices involving Hermitian conjugation. Similar modifications caused by the above extensions can be implemented in other algorithms such as addition, determinant calculation, matrix inversion, linear system solving. Such modifications will be considered in a different paper.

### v4

# Controlled measurement, Hermitian conjugation and normalization in matrix-manipulation algorithms

@@ Abstract
In this paper, we solve three important problems that are revealed, in particular, in matrix-manipulation algorithms. The principal novelty is introducing the concept of controlled measurement that solves the problem of small access probability to the desired state of ancilla and possesses several remarkable properties. We also introduce separate encoding of the real and imaginary parts of a complex matrix that allows to include the Hermitian conjugation into the list of matrix manipulations. Finally, we weaken the constraints on the modulus of matrix elements unavoidably imposed by the normalization condition for a pure quantum state. The controlled measurement together with both other extensions are implemented into the matrix multiplication algorithm. The appropriate circuits are presented.
@@ I Introduction
Quantum algorithms represent a rapidly developing area of quantum informatics whose progress is stimulated by prospective enormous advantages of quantum algorithms over their classical counterparts. We have to acknowledge such famous algorithms as the Deutsch algorithm Deu demonstrating effect of quantum parallelism in operation with Boolean functions, Shor’s factoring algorithm Sho1 ; Sho2 utilising the quantum Fourier transform QFT1 ; QFT2 ; QFT3 , Grover’s search algorithm Gro , phase-estimation algorithm for the Hermitian QFT3 and non-Hermitian Wang matrices .
@@
We consider a particular application area of quantum algorithms which covers matrix algebraic operations, such as multiplication, addition, determinant calculation, matrix inversion and its application to linear system solver. During recent years, there are many quantum algorithms solving various aspects of linear algebra. Among them are HHL-algorithm for solving systems of linear equations HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 which was applied, in particular, to Gaussian process regression ZhaoZ , algorithm for solving linear differential equations Berry , algorithm for evaluation of matrix functions Tak .
@@
A possible approach to quantum algorithms for matrix manipulations is proposed in ZZRF . This approach refers to matrices as operators to be applied to some vector. Therefore, it is essential in that approach that arbitrary nonunitary and non-Hermitian matrix must be transferred to the unitary one before been included into the quantum algorithm. Thus, first of all, one has to replace a given matrix $A$ (rectangular in general) with the Hermitian one and then transfer it to the unitary operator by exponentiating which can be approximated via Trotterization Trotter ; Suzuki ; T ; AT and Baker-Champbell-Hausdorff method Feldman ; BC . To implement such matrix transformations effectively, the set of special matrices is proposed to be prepared in advance. The phase estimation subroutine QFT3 is included into that algorithm.
@@
However, mentioned algorithms do not cover all achievement in matrix-manipulation techniques. A special block-encoding model GSLW is another approach for implementing matrix algebra. Different type of block-encoding was used in the algorithm in Ref.DFZ_2020 to embed the inverse of the matrix into the unitary transformation. In LWWZ , matrix multiplication was performed trough the binary encoding of matrix elements into the pure states of quantum systems with subsequent binary multiplication via Toffoli gates. Matrix multiplication over some rings, in particular, multiplication of Boolean matrices, was studied in KN .
@@
Some algorithms for matrix inversion have been also developed. As the first one we call the mentioned above HHL algorithm HHL where the inversion was used for solving a linear system. However, the problem of inversion of matrix eigenvalues has not been resolved in frames of quantum algorithm therein. The matrix inversion algorithm based on the singular value decomposition is presented in MRTC ; NJ . In both papers the authors apply the function evaluation problem GSLW to approximate the inverse of singular values by odd-polynomial. This approach requires introducing a special scale parameter depending on the interval where all nonzero singular values are distributed.
@@
Recently the alternative matrix-manipulation algorithms based on the encoding the matrix elements into the complex probability amplitudes of a pure superposition state of a certain subsystem have been developed QZKW_arxive2022 ; ZQKW_2024 ; ZBQKW_arXive2024 . The principal difference of such approach is that a matrix is considered as a part of the quantum state of a system rather than a unitary operator applied to this system. Therefore the elements of the resulting matrix (obtained after addition or multiplication of other matrices, or after matrix inversion) can be measured up to some constant factor. The algorithms for matrix manipulations, such as inner product, sum and product of two matrices, calculation of determinant, matrix inversion, and solving linear systems, based on special unitary operators applied to the quantum system was proposed in QZKW_arxive2022 . The realization of those unitary operators in terms of the basic quantum operations was explored later in ZQKW_2024 (inner product, matrix addition and multiplication) and then in ZBQKW_arXive2024 (calculating determinant and inverse matrix, linear system solver). Those unitary operators are based on the multi-qubit controlled operators and are supplemented by the one-qubit ancilla measurement at the final stage of the algorithm for selecting the useful result and removing the garbage. The major disadvantage of those algorithms is the small probability of obtaining the required ancilla state after measurement (small success probability). This probability decreases exponentially with the dimension of the considered matrix thus reducing efficiency of the matrix-manipulation algorithms.
The principal novelty of our paper is that we replace the above measurement of the one-qubit ancilla state with the so-called controlled measurement, i.e., the measurement controlled by the quantum state of another one-qubit ancilla. In other words, we replace the one-qubit ancilla with the two-qubit one so that the state of the first qubit governs the measurement operator applied to the second qubit. This formally simple but crucial modification removes the above disadvantage of small success probability. Unlike the algorithm implementing the usual measurement, which requires multiple running to reach the desired ancilla state, the algorithm involving the controlled measurement assumes the single run to achieve the goal. This is the main advantage of the new algorithm. Such controlled measurement can be included into all the algorithms in Refs.ZQKW_2024 ; ZBQKW_arXive2024 as a subroutine replacing the usual ancilla measurement and thus significantly increasing their relevance. We note that the controlled measurement has been implemented in Ref. ZBQKW_arXive2024 without full description. In our paper, we present the detailed analysis of this concept. The phenomenon of quantum control of classical operations (measurement), i.e., the control by the quantum state of the ancilla qubit, is a consequence of the deep relation between the quantum and classical physics and must exist relying on reality of quantum superposition states. The demonstrated advantage of the controlled measurement over the usual measurement motivates research on its particular realization in terms of acknowledged quantum and classical operators. The applicability of controlled measurement can be extended to all measurement based quantum algorithms where the success probability is of principal meaning. This refers, to the measurement-based quantum algorithms WWLN ; ZDB ; ZBD ; Wei ; WeiBook , in particular, to the HHL-algorithm for solving linear systems HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 , various applications of measurement in quantum algorithms are reviewed in USBGSN .
Along with implementing the controlled measurement, we also introduce two following extensions to the matrix-manipulation algorithms. (i) Although those algorithms deal with complex matrices in general, some operations (including Hermitian conjugation) can not be realized in frames of those settings and therefore certain constraints are to be imposed on the variety of matrix manipulations. To remove this disadvantage and construct more flexible algorithms for matrix algebra we propose to separate the real and imaginary parts of a matrix by encoding them into two orthogonal subspaces of the whole state space. This separation can be done by involving additional one-qubit subsystem whose state serves to label the real and imaginary parts of matrix elements. As a consequence, all probability amplitudes are real in such representation. We notice that the Hermitian conjugation is not a unitary operator (similar to other matrix-manipulations in ZQKW_2024 ; ZBQKW_arXive2024 ), whose realization becomes possible due to its application to the matrix encoded into the quantum state in a specific way. (ii) Another issue addressed in this paper is weakening the normalization constraint for the elements of the matrices encoded into quantum states. We recall that such constraint is imposed by the normalization condition for a pure quantum state. For that purpose we introduce an additional probability amplitude into the superposition state encoding considered matrix. This can be done involving one more auxiliary one-qubit subsystem.
Both of the above modifications are applicable to the algorithms involving encoding of particular quantum states. For instance, the already mentioned HHL-algorithm HHL uses encoding the known right-hand side of the linear system, the matrix-manipulation algorithms in ZZRF encode the vectors subjected to the matrix operations, the algorithms in ZQKW_2024 ; ZBQKW_arXive2024 encode the matrices for calculating addition, multiplication, inversion, quantum machine learning BWPRWL ; A ; RML requires encoding the input data, encoding the large data sets is utilized in the least-square linear-regression algorithms WBL ; SSP ; Wang2 and so on.
@@
Of course, the controlled measurement and both above extensions require modifications of the mentioned algorithms, in particular, the algorithms proposed in ZQKW_2024 ; ZBQKW_arXive2024 . Below we present the detailed analysis of the appropriate modifications introduced into the multiplication algorithm (see appendix in ZBQKW_arXive2024 ) and demonstrate manipulations with the input matrices involving Hermitian conjugation. However, similar modifications can be implemented in other matrix-manipulation algorithms such as inner product, addition, determinant calculation, matrix inversion, linear system solving. All quantum circuits associated with controlled measurement and two above extensions, as well as the circuit for modified multiplication algorithm, are included.